40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 392     Entry time: Mon Aug 10 15:53:46 2020
Author: Koji 
Type: General 
Category: General 
Subject: Lab status check 

Check-in to the OMC lab to see the status. Nothing seemed changed. No bug. The HEPA is running normal. The particle level was 0.

Went into the HEPA enclosure and put a cover on the OMC. Because of the gluing template, the lid could not be close completely (that's expected and fine).

The IPA vector cloth bag was not dry yet but seemed expired (some smell). There is no stock left -> 5 bags to be ordered.

ELOG V3.1.3-